OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = x + tan(x) at (pi, pi)

OpenStudy (ash2326):

@exaggerate did you try differentiating?

OpenStudy (anonymous):

yeah i got y' = 1 + sec^2(x) and when i plug in pi for x i don't know what 1 + sec^2(pi) is

OpenStudy (ash2326):

That's easy \[\sec \pi=\frac{1}{\cos\pi}\] \[cos \pi=??\]

OpenStudy (anonymous):

-sin but what about the ^2 on the secant

OpenStudy (ash2326):

\[\cos \pi= -1\] We'll use the square \[1+\sec^2 \pi=1+(-1)^2\]

OpenStudy (anonymous):

thanks i think i can get the slope now

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

i do now i just didn't know secx = 1/cosx